Please help me solve this question. Roy,Sally, and Jeff drive a total of 50 mi to work each day. Sally drives twice as far as Roy, and Jeff drives 10mi farther than Sally. Use a system of three equations in three unknowns to find how far each person drives each day.
Answer by ikleyn(52754)

.
S = 2R,              (1)
J = S + 10,          (2)
R + S + J = 50.      (3)

The setup is done. You have 3 equations in 3 unknowns.


From (2), J = S + 10 = 2R + 10,     (4)

and you can substitute ((1) and (4) into (3) to get 1 equation for 1 unknown 

R + 2R + (2R+10) = 50.


Now, I am sure, it is an easy exercise for you to complete the solution.

Good luck!